# MLKV: Multi-Layer Key-Value Heads for Memory Efficient Transformer Decoding

### v2

@@ Abstract
Auto-regressive inference of transformers benefit greatly from Key-Value (KV) caching, but can lead to major memory bottlenecks as model size, batch size, and sequence length grow at scale. We introduce Multi-Layer Key-Value (MLKV) sharing, a novel approach extending KV sharing across transformer layers to reduce memory usage beyond what was possible with Multi-Query Attention (MQA) and Grouped-Query Attention (GQA). Evaluations on various NLP benchmarks and inference metrics using uptrained Pythia-160M variants demonstrate that MLKV significantly reduces memory usage with minimal performance loss, reducing KV cache size down to a factor of 6x compared to MQA. These results highlight MLKV’s potential for efficient deployment of transformer models at scale. We provide code at https://github.com/zaydzuhri/pythia-mlkv
@@ 1 Introduction
The transformer architecture (Vaswani et al., 2017) has brought about Large Language Models (LLMs) that excel in natural language processing, but due to its auto-regressive nature when doing inference, the decoder is bottlenecked by memory bandwidth when storing and loading keys and values at each time-step, also called KV caching. Because this cache scales linearly with model size, batch size, and context length, it can even exceed the memory usage of the model weights themselves (Pope et al., 2022).
@@ 6 Discussion
The experiments show a clear trade-off between memory and accuracy. It is left to architecture designers to choose how much to sacrifice for the memory benefits, with multiple things to consider. For KV head counts above or equal to the number of layers, i.e. $mg\geq l$, given the performance shown, it is still better to use GQA/MQA instead of MLKV. We theorize that this is because having KV heads in multiple layers is more important than having multiple KV heads in the same layer. In other words, designers should sacrifice KV heads in-layer first (via GQA/MQA) and cross-layer second (via MLKV). For use cases with tighter memory requirements that require $mg<l$, then MLKV is the only way. We show that this design decision is still very much viable. We find that $mg=l/2$ with MLKV performs very near MQA, which means it should be a relatively easy decision if a KV cache is needed that is half the size of what is provided by MQA. For requirements below that, we find $mg=l/3$ and even $mg=l/6$ to still be usable, without drastic degradation. Anything below that becomes questionable. It is clear that $mg=1$ is too extreme and results in complete model collapse. The transformer benefits from the multiple recomputations of key-values from layer-to-layer, but can still compromise some level of it for the benefit of memory.
@@ 8 Conclusion
We only evaluated MLKV on decoder-only models, while encoder-decoder transformers could also benefit from this KV sharing method on their decoders. Additionally, the scale of our experiments was relatively small, conducted on models with 160 million parameters. However, models at the billion-parameter scale are becoming much more common, and MLKV has yet to be tested at such scales. Due to the relatively small size of the model we tested, the number and variety of downstream tasks that can be reliably used to compare performance is limited. Therefore, the impact of MLKV on other tasks remains to be seen. Furthermore, we did not train models from scratch and thus do not know how MLKV performs on models that are natively imbued with this KV sharing scheme from the pre-training phase.

### v3

@@ Abstract
Auto-regressive inference of transformers benefit greatly from Key-Value (KV) caching, but can lead to major memory bottlenecks as model size, batch size, and sequence length grow at scale. We introduce Multi-Layer Key-Value (MLKV) sharing, a novel approach extending KV sharing across transformer layers to reduce memory usage beyond what was possible with Multi-Query Attention (MQA) and Grouped-Query Attention (GQA). Evaluations on various NLP benchmarks and inference metrics using uptrained Pythia-160M variants demonstrate that MLKV significantly reduces memory usage with minimal performance loss, reducing KV cache size down to a factor of 6x compared to MQA. These results highlight MLKV’s potential for efficient deployment of transformer models at scale [^1] We will release all the code implementation and experimental results of our proposed architecture.
@@ 1 Introduction
The transformer architecture (Vaswani et al., 2017) has brought about Large Language Models (LLMs) that excel in natural language processing. However, due to its auto-regressive nature when doing inference, the decoder is bottlenecked by memory bandwidth when storing and loading keys and values at each time-step, also called KV caching. Because this cache scales linearly with model size, batch size, and context length, it can even exceed the memory usage of the model weights themselves (Pope et al., 2022).
@@ 5.1 Benchmarks
Model ARC-e LAMBADA PIQA SciQ Average
Pythia-410M 51.6 51.93 67.19 82.7 63.36
  Pythia-410M-MQA-24 46.04 34.06 63.82 75.1 54.76
Pythia-410M-MLKV-12 44.70 33.11 62.89 75.4 54.02
Pythia-410M-MLKV-8 46.04 31.28 62.62 72.9 53.21
Table 4: Benchmarking results of the Pythia-410M model variants. All benchmarks report accuracy and bold denotes the highest accuracy excluding the baseline Pythia-410M model.
@@ 5.2 Inference Time Measurements
Figure 5: Inference time memory measurements of the 410M parameter models. The red ’X’ indicates that beyond that batch size, an OOM error will occur.
5.3 Scaling Up
The experiments done above were done on a small scale for ablation to see what KV head counts work for optimal memory-accuracy trade-off. However, it is also important to verify whether or not these gains and losses are also applicable at a larger scale. For that, we do further uptraining and evaluation of a larger model, Pythia-410M. Specifically, since we have seen the most optimal results from configurations with $mg=l/2$ and even down to $mg=l/3$, we uptrain only these variants and also an MQA variant to compare with baseline. Since Pythia-410M has a total of 24 layers ($l=24$) and 12 query heads in each layer ($h=12$), we uptrain an MQA variant with $m=24,g=1$ resulting in Pythia-410M-MQA-24, and two MLKV variants with $m=12,g=1$ and $m=8,g=1$ resulting in Pythia-410M-MLKV-12 and Pythia-410M-MLKV-8 respectively. Results are shown in Table 4 and Figure 5.
@@ 6 Discussion
The experiments show a clear trade-off between memory and accuracy. It is left to architecture designers to choose how much to sacrifice for the memory benefits, with multiple things to consider. For KV head counts above or equal to the number of layers, i.e. $mg\geq l$, given the performance shown, it is still better to use GQA/MQA instead of MLKV. We theorize that this is because having KV heads in multiple layers is more important than having multiple KV heads in the same layer. In other words, designers should sacrifice KV heads in-layer first (via GQA/MQA) and cross-layer second (via MLKV). For use cases with tighter memory requirements that require $mg<l$, then MLKV is the only way. We show that this design decision is still very much viable. We find that $mg=l/2$ with MLKV performs very near MQA in all scales tested, which means it should be a relatively easy decision if a KV cache is needed that is half the size of what is provided by MQA. For requirements below that, we find $mg=l/3$ and even $mg=l/6$ to still be usable, without drastic degradation. Anything below that becomes questionable. It is clear that $mg=1$ is too extreme and results in model collapse. The transformer benefits from the multiple recomputations of key-values from layer-to-layer, but can still compromise some level of it for the benefit of memory.
@@ 7 Related Work
Further improvements through low-rank compression were introduced by Multi-Latent Attention (MLA) with DeepSeek-V2 (DeepSeek-AI et al., 2024). All key-values are projected down into smaller latent vectors in each layer, which are cached with a much smaller memory footprint. These vectors are then projected up into all the needed key-values for all heads in the attention mechanism. Furthermore, You-Only-Cache-Once (YOCO) (Sun et al., 2024) uses the first half of the layers to create an intermediate embedding, that is then projected into semi-global keys and values which are used for all layers of the latter half of the transformer, basically removing half of the KV cache layers.
@@ 8 Conclusion
We only evaluated MLKV on decoder-only models, while encoder-decoder transformers could also benefit from this KV sharing method on their decoders. Additionally, the scale of our experiments was relatively small, conducted on models with 160 million and 410 million parameters. However, models at the billion-parameter scale are becoming much more common, and MLKV has yet to be tested at such scales. Due to the relatively small size of the model we tested, the number and variety of downstream tasks that can be reliably used to compare performance is limited. Therefore, the impact of MLKV on other tasks remains to be seen. Furthermore, we did not train models from scratch and thus do not know how MLKV performs on models that are natively imbued with this KV sharing scheme from the pre-training phase.